OpenStudy (anonymous):

In a sample bag of M&M's candy, there are 5 brown, 6 yellow, 4 blue, 3 green, and 2 orange. What is the probability of getting 1brown and 2 orange M&M's if 3 are taken at a time? I need help working this problem out!!! please help :)

OpenStudy (anonymous):

I think there is a certain formula you have to do!

OpenStudy (anonymous):

1/228?

OpenStudy (anonymous):

0.00000128

OpenStudy (anonymous):

Can you show me how to work it out ?

OpenStudy (anonymous):

1 brown and 2 orange M&M's if 3 are taken at a time? P(1 brown and 2 orange) = [P(1 brown) + P(2 orange)]/20C3 =[P(1 brown)*P(2 orange|1 brown)]/ 20C3 =[(5/20)*2C2/19C2]/20C3 =[(1/4)*(1/171)]/1140 =[0.001462]/1140 =0.00000128

OpenStudy (anonymous):

its on this website

OpenStudy (anonymous):

That's wrong!

OpenStudy (anonymous):

Alright can you show me the correct way?

OpenStudy (raden):

yeah, that should be P(1B and 2 O's) = (5C1 * 2C2)/20C3

OpenStudy (raden):

P(1B and 2 O's) = (5C1 * 2C2)/20C3 = [5!/(5-1)!1! * 2!/(2-2)2!]/(20!/(20-3)!3!) = [5!/4!1! * 2!/0!2!]/(20!/17!3!) =(5 * 1)1140 = 5/1140 = 1/228

OpenStudy (raden):

typo, in the 3rd step =(5 * 1)/1140

OpenStudy (raden):

do u have choices as the answer, @Rea731 ?

OpenStudy (anonymous):

yes I do and that is the right answer @RadEn

OpenStudy (anonymous):

THANKS FOR THE HELP I AM PRETTY SURE I HAVE MORE MATH PROBLEMS TO COME!

OpenStudy (raden):

you're welcome if i can, i'll help :)